OpenStudy (anonymous):

Write an equation of the line through each point. Use slope-intercept form (1,-1); parallel to y=2/5x-3

OpenStudy (anonymous):

Are you having trouble with the distribution process with the 2/5?

OpenStudy (anonymous):

Eq. of any line parallel to y=2/5 x-3 is y=2/5 x+c ......(1) Because it passes through (1,-1) -1=2/5*1+c \[c=-1-\frac{ 2 }{ 5 }=\frac{ -7 }{ 5 }\] put in (1) and get the required equation.